OpenStudy (anonymous):

The price of a cellular telephone is based on peak and nonpeak service. kelsey used 45 peak and 50 nonpeak minutesand was charged $27.75. that same moth ,mitch used 70 peak minutes and 30 nonpeak minutes for a total charge of $36. what are the rates per minute for peak and nonpeak time?

OpenStudy (anonymous):

In 1990, the population of the Midwest was about 60 million. During the 1990’s, the population of this area increased an average of about 0.4 million per year. The population in the West was about 53 million in 1990. The population of this area increased an average of about 1 million per year during the 1990s.

OpenStudy (anonymous):

Assume that the rate of growth of these areas remains the same . Estimate when the population of the West Would be equal to the Midwest.

OpenStudy (anonymous):

@hartnn @mathmath333

OpenStudy (anonymous):

Sorry that i tool so long to post the question :)

OpenStudy (kamille):

hm maybe \[60(1+0.4)^{n}=53(1+1)^{n}\] where you need to find n (years) However,I dunno,maybe you had to convert 0.4 and 1 into %

OpenStudy (kamille):

it is because \[S=S _{0}(1+\frac{ p }{ 100 })^{n}\] and both of them must be equal

OpenStudy (kamille):

@amistre64

OpenStudy (anonymous):

This is what I came up with 60+.4t=p 53+t=p

OpenStudy (campbell_st):

well they look linear to me let t = years P = population (millions) MidWest: P = 0.4t + 60 West: P = t + 60 hope it makes sense

OpenStudy (campbell_st):

so all you need to do now is equate them ans solve for t 0.4t + 60 = t + 53 hope it helps

OpenStudy (anonymous):

So the answer would be 11.7 by solving those equations right ?

OpenStudy (campbell_st):

oops west should be P = t + 53

OpenStudy (campbell_st):

that's correct

OpenStudy (anonymous):

Can you help me with another?

OpenStudy (campbell_st):

if you post a new question there are lots of people who can help.

OpenStudy (anonymous):

the price of a cellular telephone is based on peak and nonpeak service. kelsey used 45 peak and 50 nonpeak minutesand was charged $27.75. that same moth ,mitch used 70 peak minutes and 30 nonpeak minutes for a total charge of $36. what are the rates per minute for peak and nonpeak time?

OpenStudy (campbell_st):

well let p = peak and n = non peak Kelsey: 45p + 50n = 27.75 Mitch: 70p + 30n = 36.00 you need to solve simultaneously I'd probably use substitution take the 1st equation and 50n = 27.75 - 45p then divide every term by 5 10n = 5.55 - 9p now substitute into the 2nd equation 70p + 3(5.55 - 9p) = 36.00 and solve for p. then when you get that you can solve for n hope it helps

OpenStudy (anonymous):

Thank you all for your help :)